# When Context Is Not Enough: Modeling Unexplained Variability in Car-Following Behavior

### v1

@@ Abstract
  Keywords:
car-following behavior
stochastic simulation
temporal correlation
Gaussian processes
interpretability
@@ 1 Introduction
Modeling and simulating car-following behaviors is a cornerstone of microscopic transportation studies and practices, directly influencing the accuracy of traffic simulations and the efficiency of transportation system control and optimization. Traditional car-following models, such as the Intelligent Driver Model (IDM) (Treiber et al.,, 2000) and Gipps’ model (Gipps,, 1981), rely on deterministic formulations to describe interactions between leading and following vehicles. While these models have proven effective in explaining the underlying mechanisms of driving behavior under various conditions, they typically treat human behavior as deterministic or incorporate only simplistic, uncorrelated noise. Such assumptions neglect the fact that driver responses are not only influenced by observable vehicle states but also shaped by latent factors, memory effects, and perception uncertainties that evolve over time (Wang et al.,, 2019). Even when noise is added, its lack of contextual awareness and temporal structure often results in simulated trajectories that miss the nuanced variability and unpredictability seen in real-world data. Importantly, a growing body of evidence suggests that human driving exhibits residual stochasticity that cannot be fully explained by context alone, necessitating a shift toward probabilistic models that explicitly account for both interpretable and unexplained sources of uncertainty.
@@
Recent studies have introduced various stochastic car-following models that explicitly incorporate randomness to capture the variability in human driving behavior. For instance, the stochastic extensions to deterministic models (Treiber et al.,, 2006; Zhang et al., 2024b, ) utilize processes like Ornstein–Uhlenbeck (OU) or autoregressive (AR) models to model fluctuations in acceleration, speed, and spacing. These approaches introduce stochastic noise terms into otherwise deterministic dynamics, enabling the representation of uncertainty in human drivers’ responses. However, because these methods typically rely on stationary or linear assumptions, their ability to capture temporal dynamics may be limited, especially in complex driving scenarios where behavioral patterns are nonlinear, context-dependent, or rapidly evolving. For example, drivers may exhibit delayed or anticipatory responses, variable reaction intensities depending on traffic states, or adaptive behaviors that change abruptly during lane changes or sudden braking events. These dynamics give rise to residual stochasticity that reflects not only context but also unobserved intent, perception errors, and memory effects—factors that traditional stochastic models generally overlook (Zhang and Sun,, 2024). Motivated by these challenges, recent work has turned to deep learning methods, particularly recurrent neural networks (RNNs) and their variants (Morton et al.,, 2016; Zhou et al.,, 2017; Wang et al.,, 2017), which have shown strong empirical performance in learning car-following patterns directly from trajectory data. These models adapt flexibly to diverse traffic contexts and behavioral styles, yet often lack explicit probabilistic structures for quantifying uncertainty or revealing the temporal dependencies in driver decision-making. As a result, it becomes difficult to interpret or diagnose why certain driving actions occur, or how prediction reliability evolves over time. This motivates the need for hybrid approaches that combine the representational power of deep models with structured components that offer interpretability and probabilistic insight.
@@
This work proposes an interpretable stochastic car-following framework that integrates deep neural sequence modeling with nonstationary GPs. Building on the DeepAR architecture (Salinas et al.,, 2020), our model predicts both the mean acceleration and a context-dependent covariance structure via a scenario-adaptive Gibbs kernel. At each time step, the recurrent network ingests spacing, speed, and relative speed to output a predicted acceleration, a lengthscale, and a variance. These kernel parameters modulate a Gaussian-process residual that captures temporally correlated deviations from the mean dynamics, allowing the model to adapt its memory and uncertainty to evolving traffic conditions in an end-to-end trainable manner.
The primary contribution of this work is the development of a stochastic modeling framework for car-following behavior that explicitly captures context-dependent temporal correlations in human driving. This core modeling innovation delivers three key benefits. First, it provides a principled way to represent human driving variability as a temporally correlated stochastic process, overcoming the stationarity assumptions of classical models. Second, the scenario-adaptive Gibbs kernel enables the model to learn dynamic correlation scales and heteroskedastic noise levels that reflect different driving regimes in a data-driven and interpretable manner. Finally, extensive experiments on naturalistic HighD trajectories show that our approach yields lower simulation errors and sharper, better-calibrated uncertainty estimates than both deterministic baselines and stationary GP variants. Together, these advantages make the framework a powerful and interpretable tool for high-fidelity traffic simulation and safety-critical behavior analysis.
@@
2 Related Works
Understanding and modeling car-following behavior has been a longstanding focus in transportation research. Over the decades, a wide range of models have been proposed, from rule-based formulations grounded in traffic flow theory to flexible machine learning approaches trained on naturalistic driving data. Comprehensive surveys such as Chen et al., 2023b and Zhang et al., 2024c provide detailed overviews of this evolving landscape. Building on this foundation, this section highlights two major lines of research most relevant to our study: (i) traditional analytical models, typically deterministic in nature, augmented with stochastic components to capture behavioral variability, and (ii) data-driven methods, particularly those leveraging deep learning and sequential modeling to learn complex driving patterns directly from observed trajectories.
@@ 2.1 Traditional Analytical Models with Stochastic Extensions
Traditional car-following models such as the General Motors (GM) stimulus–response model, the Optimal Velocity Model (OVM), and the IDM have played a foundational role in the development of microscopic traffic theories. The GM model links a follower vehicle’s acceleration to the relative speed and headway with respect to the leader, providing an intuitive yet simplistic response mechanism (Gazis et al.,, 1961). The OVM introduces a desired speed function that depends on headway, capturing emergent traffic phenomena such as spontaneous congestion waves and phase transitions between free-flow and stop-and-go conditions (Bando et al.,, 1995). The IDM advances this framework by incorporating a driver’s desired speed, time gap, and acceleration limits into a smooth and collision-avoiding formulation (Treiber et al.,, 2000). While these deterministic models offer analytical tractability and interpretability, they rely on the assumption of homogeneity across drivers and fixed parameter values. As a result, they often fail to replicate the variability observed in empirical driving data, particularly under conditions involving abrupt maneuvers, inter-driver differences, or transient traffic regimes. This has motivated the exploration of probabilistic modeling paradigms that account for uncertainty and behavioral heterogeneity.
@@
In response, Bayesian methods have been widely adopted to enrich classical models with stochastic structure. By treating model parameters as random variables, Bayesian calibration enables the quantification of uncertainty and facilitates more robust fitting to observed trajectories. For instance, recent studies employ Bayesian dynamic regression and Markov Chain Monte Carlo (MCMC) techniques to infer posterior distributions over IDM parameters, thereby capturing inter-driver variability and enhancing predictive realism (Zhang et al., 2024b, ). Complementary approaches such as Approximate Bayesian Computation (ABC) have also shown promise, especially in scenarios where likelihood functions are intractable or unknown (Jiang et al.,, 2025).
@@
Beyond parameter uncertainty, a parallel line of research has leveraged latent-variable models to capture the multi-modality and nonstationarity of driving behavior. Hidden Markov Models (HMMs) and Gaussian Mixture Models (GMMs) are frequently used to represent discrete behavioral modes, such as accelerating, cruising, and braking, and to infer regime-switching dynamics from trajectory data (Zhang et al.,, 2025; Vaitkus et al.,, 2014). These models are particularly effective at reproducing realistic stop-and-go patterns and behavioral transitions. Related work on mixture models and latent class frameworks further extends this idea by probabilistically clustering driver types or behavioral styles (Chen et al., 2023a, ; Zhang et al., 2024a, ), offering a principled way to represent population-level heterogeneity.
@@
More recently, GP models have emerged as a non-parametric Bayesian alternative for learning car-following dynamics directly from data. By modeling acceleration as a stochastic function of spacing, speed, and relative speed, GPs provide flexible representations with calibrated confidence intervals. Their ability to capture smooth, data-driven response surfaces makes them well-suited for modeling uncertainty in driver responses, although their use in traffic modeling has only recently gained traction (Soldevila et al.,, 2021; Wang et al.,, 2022).
@@ 2.2 Data-Driven and Neural Sequence Modeling Approaches
In recent years, the development of large-scale trajectory datasets and advances in deep learning have fueled growing interest in using neural networks to model driving behavior directly from data (Li et al.,, 2020). Unlike traditional models that rely on manually crafted rules, data-driven approaches can learn complex, non-linear response patterns by optimizing over observed driving trajectories (He et al.,, 2015; Huang et al.,, 2025). Among these, recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) architectures, have proven especially effective in capturing temporal dependencies inherent in car-following dynamics.
@@
For instance, Zhou et al., (2017) introduced a basic RNN-based car-following model that successfully reproduced traffic oscillation patterns within vehicle platoons. Building on this, LSTM-based models have shown enhanced capabilities in learning long-range temporal correlations by mitigating issues of vanishing gradients, making them suitable for representing delayed driver reactions and cumulative behavioral patterns (Ma and Qu,, 2020). Empirical studies have consistently demonstrated that LSTM- and GRU-based models outperform classical car-following models in reproducing acceleration and braking behaviors, particularly in complex or congested traffic scenarios (Wang et al.,, 2017; Lin et al.,, 2020). A further extension of these ideas is found in probabilistic sequence models, such as DeepAR (Salinas et al.,, 2020), which formulate driving behavior forecasting as an autoregressive sequence prediction problem. Rather than providing point estimates alone, DeepAR learns a full predictive distribution over future trajectories, thus offering both forecasts and associated uncertainty intervals. When trained on car-following sequences, these models can generate calibrated probabilistic forecasts that are crucial for risk-aware decision-making in autonomous driving systems. The ability to model uncertainty makes these approaches particularly well-suited for the inherently stochastic and context-sensitive nature of human driving. Beyond recurrent models, several studies have benchmarked data-driven architectures, including feedforward neural networks, LSTMs, and reinforcement learning agents, against traditional physics-based models (Huang et al.,, 2021). Results suggest that data-driven models can more accurately reproduce observed vehicle trajectories while better capturing nuanced behavioral patterns that are often missed by rule-based formulations (Zhu et al.,, 2018). These findings highlight the promise of modern neural architectures in learning rich representations of driving dynamics directly from data, thereby reducing reliance on hand-engineered assumptions.
@@
More recently, hybrid modeling frameworks have emerged as a promising direction, aiming to bridge the gap between physical interpretability and data-driven flexibility. Rather than discarding classical models, these approaches integrate them into neural architectures as structured priors or corrective baselines. For example, Wang and Feng, (2022) proposed a neural-augmented IDM framework, where a neural network learns to compensate for discrepancies between the IDM and empirical behavior. Similarly, Soldevila et al., (2021) combined a parametric car-following model with a non-parametric GP module, allowing the model to learn data-driven corrections while preserving the theoretical structure of the base model. An interesting instance of hybridization is the use of time-varying parameters governed by latent embeddings, as demonstrated in Zhang et al., (2022), where the IDM is conditioned on a learned driver-style vector inferred through a neural process. This design allows the model to generalize across a range of behavioral regimes while maintaining interpretability. However, we note a common misconception in interpreting such frameworks: the parameters are often described as “time-varying,” when in fact, they are better understood as context-dependent. That is, the variation in parameters arises from the evolving driving state, including spacing, speed, and relative speed, rather than from an explicit dependence on time itself. Misinterpreting context-driven adaptation as purely temporal variation risks obscuring the structured and reactive nature of human driving behavior, and may lead to incorrect assumptions about model design and interpretation. By embedding stochastic processes such as GPs or neural processes within deep architectures, these hybrid models achieve a balance between generalization and data efficiency. Grounding the learning process in well-established physical principles not only enhances interpretability but also helps ensure safety and constraint adherence, which is the critical aspects in traffic simulation and autonomous driving applications.
@@ 3.1 Problem Formulation
Traditional car-following models, such as the IDM (Treiber et al.,, 2000), represent $f_{\mathrm{CF}}(\cdot)$ analytically. These approaches simplify the complexity of driving behaviors by assuming static functional relationships between the covariates. However, human driving behaviors exhibit significant temporal variability and uncertainty, which deterministic models often fail to capture. In the following, we aim to enhance the modeling of car-following behaviors by leveraging sequential data $\{\bm{s},\Delta\bm{v},\bm{v}\}$ to account for the inherent stochasticity and temporal dependencies in driving behaviors.

### v2

@@ 1 Introduction
Modeling and simulating car-following behaviors is a cornerstone of microscopic transportation studies and practice, directly influencing the fidelity of traffic simulation and the effectiveness of transportation control and optimization. Traditional car-following models, such as the Intelligent Driver Model (IDM) (Treiber et al., 2000) and Gipps’ model (Gipps, 1981), typically use deterministic rules to represent interactions between leading and following vehicles. While these models capture key behavioral mechanisms, they often treat human responses as deterministic or introduce only simplistic, temporally uncorrelated disturbances. Such assumptions are limiting because driver actions are shaped not only by observable kinematic states but also by unobserved and time-varying factors, including memory and adaptation, perception uncertainty, and latent situational cues (e.g., brake-light signals, weather and visibility, road geometry, and driver internal states) that are rarely available in standard trajectory datasets (Wang et al., 2019). Consequently, even when conditioning on common car-following stimuli (e.g., spacing and relative speed), the remaining discrepancy typically exhibits persistent stochastic components and regime-dependent variability rather than behaving as i.i.d. noise. This view is supported by trajectory-data and experimental evidence and has motivated stochastic car-following formulations that model acceleration or speed fluctuations as random processes, demonstrating that structured stochasticity can be essential for reproducing observed oscillations and variability in traffic dynamics (Laval et al., 2014; Ngoduy et al., 2019; Xu and Laval, 2020).
@@
Recent studies have introduced various stochastic car-following models that explicitly incorporate randomness to capture the variability in human driving behavior. For instance, the stochastic extensions to deterministic models (Treiber et al., 2006; Zhang et al., 2024b) utilize processes like Ornstein–Uhlenbeck (OU) or autoregressive (AR) models to model fluctuations in acceleration, speed, and spacing. These approaches introduce stochastic noise terms into otherwise deterministic dynamics, enabling the representation of uncertainty in human drivers’ responses. However, because these methods typically rely on stationary or linear assumptions, their ability to capture temporal dynamics may be limited, especially in complex driving scenarios where behavioral patterns are nonlinear, context-dependent, or rapidly evolving. For example, drivers may exhibit delayed or anticipatory responses, variable reaction intensities depending on traffic states, or adaptive behaviors that change abruptly during lane changes or sudden braking events. These dynamics give rise to residual stochasticity that reflects not only context but also unobserved intent, perception errors, and memory effects—factors that traditional stochastic models generally overlook (Zhang and Sun, 2024; Treiber et al., 2006; Bifulco et al., 2013; Wang et al., 2020). Motivated by these challenges, recent work has turned to deep learning methods, particularly recurrent neural networks (RNNs) and their variants (Morton et al., 2016; Zhou et al., 2017; Wang et al., 2017), which have shown strong empirical performance in learning car-following patterns directly from trajectory data. These models adapt flexibly to diverse traffic contexts and behavioral styles, yet often lack explicit probabilistic structures for quantifying uncertainty or revealing the temporal dependencies in driver decision-making. As a result, it becomes difficult to interpret or diagnose why certain driving actions occur, or how prediction reliability evolves over time. This motivates the need for hybrid approaches that combine the representational power of deep models with structured components that offer interpretability and probabilistic insight.
@@
This work proposes an interpretable stochastic car-following framework that integrates deep neural sequence modeling with nonstationary GPs. Using DeepAR (Salinas et al., 2020) as one instantiation of the mean dynamics, our model predicts both the mean acceleration and a context-dependent covariance structure via a scenario-adaptive Gibbs kernel. At each time step, the recurrent network ingests spacing, speed, and relative speed to output a predicted acceleration, a lengthscale, and a variance. These kernel parameters modulate a Gaussian-process residual that captures temporally correlated deviations from the mean dynamics, allowing the model to adapt its memory and uncertainty to evolving traffic conditions in an end-to-end trainable manner.
The primary contribution of this work is the development of a stochastic modeling framework for car-following behavior that explicitly captures context-dependent temporal correlations in human driving. Conceptually, we distinguish aleatoric uncertainty arising from irreducible variability in driving actions and sensing/measurement noise from epistemic uncertainty due to limited coverage of driving contexts or model misspecification. Our framework focuses on representing structured aleatoric variability through a temporally correlated residual process whose correlation and variance adapt to driving context, while providing a clear lens for interpreting potential epistemic effects under domain shift or sparse contexts. This core modeling innovation delivers three key benefits. First, it provides a principled way to represent human driving variability as a temporally correlated stochastic process, overcoming the stationarity assumptions of classical models. Second, the scenario-adaptive Gibbs kernel enables the model to learn dynamic correlation scales and heteroskedastic noise levels that reflect different driving regimes in a data-driven and interpretable manner. Finally, extensive experiments on naturalistic HighD trajectories show that our approach yields lower simulation errors and sharper, better-calibrated uncertainty estimates than both deterministic baselines and stationary GP variants. Together, these advantages make the framework a powerful and interpretable tool for high-fidelity traffic simulation and safety-critical behavior analysis.
@@
2 Related Work
Understanding and modeling car-following behavior has been a longstanding focus in transportation research. Over the decades, a wide range of models have been proposed, from rule-based formulations grounded in traffic flow theory to flexible machine learning approaches trained on naturalistic driving data (Zheng et al., 2023; He et al., 2022). Comprehensive surveys such as Chen et al. (2023a) and Zhang et al. (2024c) provide detailed overviews of this evolving landscape. Building on this foundation, this section highlights two major lines of research most relevant to our study: (i) traditional analytical models, typically deterministic in nature, augmented with stochastic components to capture behavioral variability, and (ii) data-driven methods, particularly those leveraging deep learning and sequential modeling to learn complex driving patterns directly from observed trajectories.
@@ 2.1 Traditional Analytical Models with Stochastic Extensions
Traditional car-following models such as the General Motors (GM) stimulus–response model, the Optimal Velocity Model (OVM), and the IDM have played a foundational role in the development of microscopic traffic theories. The GM model links a follower vehicle’s acceleration to the relative speed and headway with respect to the leader, providing an intuitive yet simplistic response mechanism (Gazis et al., 1961). The OVM introduces a desired speed function that depends on headway, capturing emergent traffic phenomena such as spontaneous congestion waves and phase transitions between free-flow and stop-and-go conditions (Bando et al., 1995). The IDM advances this framework by incorporating a driver’s desired speed, time gap, and acceleration limits into a smooth and collision-avoiding formulation (Treiber et al., 2000). Beyond traditional stimulus-response frameworks, more sophisticated analytical models have been developed to account for human physiological and psychological constraints. For instance, psychophysical models, such as the Wiedemann model (Wiedemann, 1974), utilize action-point thresholds to describe discontinuities in acceleration, reflecting the limits of human perception in detecting small changes in relative speed or distance. While driver heterogeneity can be incorporated in simulation practice (e.g., via driver classes or parameter distributions), many analytical car-following models are still often deployed with fixed or slowly varying parameterizations within a given run and with simplified disturbance assumptions (frequently i.i.d. and context-invariant). Even when multiple parameter sets are used across traffic states, the remaining fluctuations are rarely modeled as a context-dependent, temporally correlated residual process. This motivates our framework, which complements existing mean-model formulations by learning regime-dependent variance and temporal correlation in the residual and can be integrated on top of heterogeneous parameterizations.
@@
In response, Bayesian methods have been widely adopted to enrich classical models with stochastic structure. By treating model parameters as random variables, Bayesian calibration enables the quantification of uncertainty and facilitates more robust fitting to observed trajectories. For instance, recent studies employ Bayesian dynamic regression and Markov Chain Monte Carlo (MCMC) techniques to infer posterior distributions over IDM parameters, thereby capturing inter-driver variability and enhancing predictive realism (Zhang et al., 2024b). Complementary approaches such as Approximate Bayesian Computation (ABC) have also shown promise, especially in scenarios where likelihood functions are intractable or unknown (Jiang et al., 2025).
@@
Beyond parameter uncertainty, a parallel line of research has leveraged latent-variable models to capture the multi-modality and nonstationarity of driving behavior. Hidden Markov Models (HMMs) and Gaussian Mixture Models (GMMs) are frequently used to represent discrete behavioral modes, such as accelerating, cruising, and braking, and to infer regime-switching dynamics from trajectory data (Zhang et al., 2025; Vaitkus et al., 2014). These models are particularly effective at reproducing realistic stop-and-go patterns and behavioral transitions. Related work on mixture models and latent class frameworks further extends this idea by probabilistically clustering driver types or behavioral styles (Chen et al., 2023b; Zhang et al., 2024a), offering a principled way to represent population-level heterogeneity.
@@
More recently, GP models have emerged as a non-parametric Bayesian alternative for learning car-following dynamics directly from data. By modeling acceleration as a stochastic function of spacing, speed, and relative speed, GPs provide flexible representations with calibrated confidence intervals. Their ability to capture smooth, data-driven response surfaces makes them well-suited for modeling uncertainty in driver responses, although their use in traffic modeling has only recently gained traction (Soldevila et al., 2021; Wang et al., 2022).
@@ 2.2 Data-Driven and Neural Sequence Modeling Approaches
In recent years, the development of large-scale trajectory datasets and advances in deep learning have fueled growing interest in using neural networks to model driving behavior directly from data (Li et al., 2020). Unlike traditional models that rely on manually crafted rules, data-driven approaches can learn complex, non-linear response patterns by optimizing over observed driving trajectories (He et al., 2015; Huang et al., 2025). Among these, recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) architectures, have proven especially effective in capturing temporal dependencies inherent in car-following dynamics.
@@
For instance, Zhou et al. (2017) introduced a basic RNN-based car-following model that successfully reproduced traffic oscillation patterns within vehicle platoons. Building on this, LSTM-based models have shown enhanced capabilities in learning long-range temporal correlations by mitigating issues of vanishing gradients, making them suitable for representing delayed driver reactions and cumulative behavioral patterns (Ma and Qu, 2020). Empirical studies have consistently demonstrated that LSTM- and GRU-based models outperform classical car-following models in reproducing acceleration and braking behaviors, particularly in complex or congested traffic scenarios (Wang et al., 2017; Lin et al., 2020). A further extension of these ideas is found in probabilistic sequence models, such as DeepAR (Salinas et al., 2020), which formulate driving behavior forecasting as an autoregressive sequence prediction problem. Rather than providing point estimates alone, DeepAR learns a full predictive distribution over future trajectories, thus offering both forecasts and associated uncertainty intervals. When trained on car-following sequences, these models can generate calibrated probabilistic forecasts that are crucial for risk-aware decision-making in autonomous driving systems. The ability to model uncertainty makes these approaches particularly well-suited for the inherently stochastic and context-sensitive nature of human driving. Beyond recurrent models, several studies have benchmarked data-driven architectures, including feedforward neural networks, LSTMs, and reinforcement learning agents, against traditional physics-based models (Huang et al., 2021). Results suggest that data-driven models can more accurately reproduce observed vehicle trajectories while better capturing nuanced behavioral patterns that are often missed by rule-based formulations (Zhu et al., 2018). These findings highlight the promise of modern neural architectures in learning rich representations of driving dynamics directly from data, thereby reducing reliance on hand-engineered assumptions.
@@
More recently, hybrid modeling frameworks have emerged as a promising direction, aiming to bridge the gap between physical interpretability and data-driven flexibility. Rather than discarding classical models, these approaches integrate them into neural architectures as structured priors or corrective baselines. For example, Wang and Feng (2024) proposed a neural-augmented IDM framework, where a neural network learns to compensate for discrepancies between the IDM and empirical behavior. Similarly, Soldevila et al. (2021) combined a parametric car-following model with a non-parametric GP module, allowing the model to learn data-driven corrections while preserving the theoretical structure of the base model. An interesting instance of hybridization is the use of time-varying parameters governed by latent embeddings, as demonstrated in Zhang et al. (2022), where the IDM is conditioned on a learned driver-style vector inferred through a neural process. This design allows the model to generalize across a range of behavioral regimes while maintaining interpretability. However, we note a common misconception in interpreting such frameworks: the parameters are often described as “time-varying,” when in fact, they are better understood as context-dependent. That is, the variation in parameters arises from the evolving driving state, including spacing, speed, and relative speed, rather than from an explicit dependence on time itself. Misinterpreting context-driven adaptation as purely temporal variation risks obscuring the structured and reactive nature of human driving behavior, and may lead to incorrect assumptions about model design and interpretation. By embedding stochastic processes such as GPs or neural processes within deep architectures, these hybrid models achieve a balance between generalization and data efficiency. Grounding the learning process in well-established physical principles not only enhances interpretability but also helps ensure safety and constraint adherence, which is the critical aspects in traffic simulation and autonomous driving applications.
@@ 3.1 Problem Formulation
Traditional car-following models, such as the IDM (Treiber et al., 2000), represent $f_{\mathrm{CF}}(\cdot)$ analytically. These approaches simplify the complexity of driving behaviors by assuming static functional relationships between the covariates. However, human driving behaviors exhibit significant temporal variability and uncertainty, which deterministic models often fail to capture. In the following, we aim to enhance the modeling of car-following behaviors by leveraging sequential data $\{\bm{s},\Delta\bm{v},\bm{v}\}$ to account for the inherent stochasticity and temporal dependencies in driving behaviors.